OpenStudy (anonymous):

What is the value of x?: (11^-4)^-6=11^x

OpenStudy (irishboy123):

\[(11^{-4})^{-6}=11^x\] https://gyazo.com/0581545a3ffd253ccb4b631ac03d98db

OpenStudy (anonymous):

nope didn't work

OpenStudy (dinamix):

x=24

OpenStudy (dinamix):

@leeeeeevi1

OpenStudy (anonymous):

Thanks dinamix

OpenStudy (dinamix):

\[24\ln11 = x \ln 11 \] \[24 = x\] this method @leeeeeevi1

OpenStudy (dinamix):

@IrishBoy123

OpenStudy (irishboy123):

it is important to know that \[\large (11^{−4})^{−6}=11^{(-4 \ \times -6)}\] that is what the link tells us https://gyazo.com/4b5d2b0e07a7b22f9b70f72f6c9bd7b6